Shift summary, 07:00 - 15:30, Tuesday June 1st.

Bruce Winstein (bruce@hep.uchicago.edu)
Tue, 1 Jun 1999 15:52:52 -0500

Shift Summary for Day shift, 06/01/99

On Shift:
Vivian O'Dell
Bruce Winstein

Other People Present: lots (see below)

Hours of beam delivered at all intensities: 0.0
Hours of beam delivered at nominal intensity: 0.0
Hours of E832 data (special runs not included): 0.0
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time:
waiting for Beam.

========================================================================
Runs ended this shift:

Tapes brought to FCC this shift: 0

========================================================================
Accesses during shift: various fixes (see below)

========================================================================
Description of Shift:

Main Injector dedication day -> no beam in the Main Injector
Scripts for setting magnets work; documentation in place
Earl
Continuing efforts to work on PT kick problem from owl shift
Rick, Ed, and others by telephone
Poor looking chamber histos from last night's hadron running still
under study
Sasha, Hogan
Difference in monitoring vs. offline histograms traced to
different trigger mixtures
Sasha
DC ramping signals modified to start(end) on T4(6)
Miguel
Nim bin for CsI HV interlock replaced
Ron, Julie
Remote recycling of AC power for laser repeater installed
Bob T., Miguel